# Is JPEG AI going to change image forensics?

### v1

@@ 9 Image splicing localization additional results
Visual inspection. Fig. 12 to Fig. 15 report some examples of output maps for both the TruFor and MMFusion detectors. In all plots, the detectors process the same DSO1 input image compressed with different JPEG AI ratios.
@@
However, it is interesting to notice that, for certain samples, increasing the BPP during compression does not lead to a better . For example, the output of MMFusion in Fig. 14 maintains a high number of false positives even for and , and TruFor shows a similar behavior in Fig. 15. Therefore, while the overall tendency highlighted in Tab. 5 is clear, these examples seem to suggest that there is no guarantee that compressing images with high BPP values will not interfere with the detectors responses.
Quantitative analysis. Tab. 24 and Tab. 25 report the quantitative analysis results for the two detectors, showing a breakdown for each dataset. As summarized in Tab. 5 of the main paper, we can see a decrease in localization performances for all datasets and both detectors as we increase the ratio of the JPEG AI compression.
From this perspective, it is interesting to analyze the CASIA [34] results. CASIA samples are natively compressed in JPEG format. We can see that both detectors have huge performances in localizing the manipulation on original CASIA images, therefore indicating a certain robustness to standard JPEG. However, performances once again drop whenever we compress these images with JPEG AI.
To verify this intuition, i.e., that JPEG compression is not as dangerous as JPEG AI, we perform an additional test by compressing with JPEG at varying QFs one of the datasets natively provided in uncompressed format. Tab. 26 reports the splicing localization results of TruFor on the DSO1 dataset compressed with JPEG at varying QF values. While we can still notice a certain degradation in performances, especially at low QF, numbers are still very high compared to those obtained by compressing the samples with JPEG AI.
| Metric | Original | JPEG AI compressed (BPP) | ||||||
| CASIA | ||||||||
| @0.5 | ||||||||
| best | ||||||||
| F1@0.5 | ||||||||
| F1 best | ||||||||
| FPR@0.5 | ||||||||
| CocoGlide | ||||||||
| @0.5 | ||||||||
| best | ||||||||
| F1@0.5 | ||||||||
| F1 best | ||||||||
| FPR@0.5 | ||||||||
| Columbia | ||||||||
| @0.5 | ||||||||
| best | ||||||||
| F1@0.5 | ||||||||
| F1 best | ||||||||
| FPR@0.5 | ||||||||
| Coverage | ||||||||
| @0.5 | ||||||||
| best | ||||||||
| F1@0.5 | ||||||||
| F1 best | ||||||||
| FPR@0.5 | ||||||||
| DSO1 | ||||||||
| @0.5 | ||||||||
| best | ||||||||
| F1@0.5 | ||||||||
| F1 best | ||||||||
| FPR@0.5 | ||||||||
| Metric | Original | JPEG AI compressed (BPP) | ||||||
| CASIA | ||||||||
| @0.5 | ||||||||
| best | ||||||||
| F1@0.5 | ||||||||
| F1 best | ||||||||
| FPR@0.5 | ||||||||
| CocoGlide | ||||||||
| @0.5 | ||||||||
| best | ||||||||
| F1@0.5 | ||||||||
| F1 best | ||||||||
| FPR@0.5 | ||||||||
| Columbia | ||||||||
| @0.5 | ||||||||
| best | ||||||||
| F1@0.5 | ||||||||
| F1 best | ||||||||
| FPR@0.5 | ||||||||
| Coverage | ||||||||
| @0.5 | ||||||||
| best | ||||||||
| F1@0.5 | ||||||||
| F1 best | ||||||||
| FPR@0.5 | ||||||||
| DSO1 | ||||||||
| @0.5 | ||||||||
| best | ||||||||
| F1@0.5 | ||||||||
| F1 best | ||||||||
| FPR@0.5 | ||||||||
| Metric | Original | JPEG Quality Factor (QF) | |||
| 65 | 75 | 85 | 95 | ||
| best | |||||
| F1@0.5 | |||||
| F1 best | |||||
| FPR @0.5 | |||||

### v2

@@ 10 Image splicing localization additional results
Fig. 14 to Fig. 17 report some examples of output maps for the TruFor and MMFusion detectors. In all plots, the detectors process the same DSO1 input image compressed with different JPEG AI BPP.
@@
However, it is interesting to notice that, for certain samples, increasing the BPP during compression does not lead to a better . For example, the output of MMFusion in Fig. 16 maintains a high number of false positives even for and , and TruFor shows a similar behavior in Fig. 17. Therefore, while the overall tendency highlighted in Tab. 5 is clear, these examples seem to suggest that there is no guarantee that compressing images with high BPP values will not interfere with the responses of the detectors.
11 Double JPEG AI compression
The study field relative to multiple JPEG compression is lively [73]. Therefore, in the following, we briefly analyze some of the implications of double JPEG AI compression. Please notice that this analysis is not exhaustive, and we plan to explore the more profound implications of double JPEG AI in future studies.
Deepfake image detection. Fig. 18 reports the scores of the [42] detector extracted from pristine images compressed in the scenarios illustrated in Sec. 5.3, namely i) single JPEG AI compression, ii) double JPEG AI compression where the original-format images have been first compressed with the lowest BPP considered, i.e., BPP , and then re-compressed at various BPP values, and iii) the original-format images have been first compressed with the highest BPP considered, i.e., BPP , and then re-compressed at various BPP values.
Fig. 18(b) reports the results for the first scenario. As we can quickly inspect, regardless of the BPP value used in the second compression, the scores distribution of pristine samples will lie above the threshold in the area corresponding to the scores distribution of single compressed BPP images. On the contrary, Fig. 18(c) shows that if we first compress the samples with a high BPP value, the distribution of the scores will move accordingly to the BPP value of the second compression operation, i.e., a BPP after a BPP compression will shift the scores in the same area as a single BPP compression. These results suggest that JPEG AI behaves similarly to standard compression formats in the sense that, if a strong compression has been applied to the input samples, the compressed images will likely still bear forensic artifacts relative to the first compression stage regardless of the BPP used in the second one.
Image splicing localization. In image splicing localization, the double compression usually does not interest the whole image. Indeed, it is rather common that only the spliced area indicated by the tampering mask presents artifacts relative to a previous compression operation. This is because the source and target samples might have been saved with different formats or the same coding scheme but with one of the two images being processed with a more aggressive configuration.
We replicated this scenario by taking a spliced sample from an uncompressed dataset, i.e., DSO-1, and compressing this sample with all the BPP values considered in our experiments. After that, we created several spliced images where the spliced area was JPEG AI compressed at various BPP, and the surrounding content was uncompressed. We finally re-compressed all these images with BPP equal to and . In this way, we are creating a mismatch in the spatial distribution of compression traces due to the spliced area bearing traces of the first JPEG AI compression, similar to the scenario analyzed in double JPEG compression.
Fig. 19 and Fig. 20 report the results of our double JPEG AI splicing operation together with the mask estimated by MMFusion. We also include the results of single-compressing the original format spliced image for reference. As we can quickly inspect, we can draw similar conclusions to the deepfake image detection experiments. In Fig. 20, the images were saved with BPP , i.e., with very high quality. In this scenario, if the spliced area has been initially compressed with low BPP, MMFusion can detect this inconsistency and localize the spliced area with results much better than the reference single BPP JPEG AI compression. On the other hand, in Fig. 19, the second compression stage at BPP likely erases all the traces relative to the first, hence making MMFusion deliver results very similar to a single JPEG AI compression. These results suggest that the inconsistency in the spatial distribution of JPEG AI artifacts might become a very powerful forensic instrument.